OpenStudy (anonymous):

What is the denominator of the simplified form of (1/2)-(3/(4x-2)) a.) 2 b.) x-2 c.) 2x-1

OpenStudy (anonymous):

i got b is that right?

OpenStudy (anonymous):

\[\frac{1}{2} - \frac{3}{4x - 2}\]

OpenStudy (anonymous):

ya thats the equation

OpenStudy (anonymous):

\[\frac{1}{2} - \frac{3}{2(2x - 1)}\] Find a common denominator: \[(\frac{1}{2} \times \frac{2x - 1}{2x - 1}) - \frac{3}{2(2x - 1)}\] \[\frac{2x - 1}{2(2x - 1)} - \frac{3}{2(2x - 1)} = \frac{2x - 1 - 3}{2(2x - 1)}\]

OpenStudy (anonymous):

Now simplify

OpenStudy (anonymous):

so the answer is c?